# Identifying Dwarfs of MC Analog GalaxiEs (ID-MAGE): The Search for Satellites Around Low-Mass Hosts

### v3

@@ Appendix A Environment Plots
This section provides details on the environment of specific host galaxies. The online figure set for Figure 2 illustrates the environment of each host along with its candidate satellites. The surveyed areas are shaded gray, and prominent galaxies are displayed with their approximate virial radii: 110 kpc for SMC-mass galaxies, 150 kpc for LMC-mass galaxies, and 300 kpc for more massive galaxies. In some cases, host search areas overlap, causing a few candidates to be found in multiple hosts’ search areas. If a candidate is found in more than one searched area, we associate it with the physically closest host, assuming it lies at the same distance. Follow-up observations (see Section 6) will confirm the true hosts of these candidates.
@@
NGC 4244: NGC 4244 (4.20 Mpc) appears to be part of a group of LMC-mass and smaller galaxies, including NGC 4395(4.65 Mpc)). This group of low-mass galaxies has measured TRGB distances between 4 and 5 Mpc (Tully et al., 2009). The galaxies CGCG187-05, UGC 7559, and UGC 7599 lie just beyond NGC 4244’s virial radius, with TRGB distances of 4.85 Mpc, 4.97 Mpc, and 4.72 Mpc, respectively (Tully et al., 2009). These distances place them more than 500 kpc behind NGC 4244, suggesting they belong to the same galaxy group but are unlikely to be its satellites.
NGC 4449: NGC 4449’s (4.16 Mpc) virial radius overlaps with that of M94, which is at a similar distance (4.3 Mpc; Tully et al. 2009), indicating that NGC 4449 and its satellites are likely members of the larger M94 galaxy group.
IC 4182: IC 4182’s (4.24 Mpc) virial radius overlaps with that of M94, which is at a similar distance (4.3 Mpc Tully et al. 2009), indicating that IC 4182 and its satellites are likely members of the larger M94 galaxy group.
NGC 4236: NGC 4236’s (4.31 Mpc) virial radius overlaps with that of UGC 07490; however, the two are unlikely to be associated due to their large difference in velocities. In Cosmicflows-4, NGC 4236 has a measured velocity of 0 km s-1 (Tully et al., 2009), while UGC 07490 has a velocity of 4685 km s-1.
@@
NGC 4395: NGC 4395 (4.65 Mpc) appears to be part of a group of LMC-mass and smaller galaxies, including NGC 4244 (4.20 Mpc). This group has measured TRGB distances between 4 and 5 Mpc (Tully et al., 2009). UGC 07698 and UGC 07605 lie just beyond NGC 4395’s estimated virial radius and appear to belong to the same galaxy group based on their Cosmicflows-4 distances.
NGC 5585: NGC 5585 (6.84 Mpc)) is located on the outskirts of a galaxy group. Its virial radius overlaps that of M101, which is at a similar distance (6.7 Mpc; Tully et al. 2009).
@@
ESO486-G21: For ESO486-G21 (9.7 Mpc), the nearby galaxy NGC 1744 does not have a TRGB distance measurement but has a measured velocity of 741 km s-1 (Springob et al., 2005). This is similar to ESO486-G21’s velocity of 840 km s-1 (Meyer et al., 2004). Additionally, multiple Tully-Fischer distance measurements place NGC 1744 within 1 Mpc of ESO486-G21 (e.g. Tully et al. 2009), suggesting that these two galaxies may be associated and part of the same group.

### v5

@@ Appendix A Environment Plots
This section provides details on the environment of specific host galaxies. The online figure set for Figure 2 illustrates the environment of each host along with its candidate satellites. The surveyed areas are shaded gray, and prominent galaxies are displayed with their approximate virial radii: 110 kpc for SMC-mass galaxies, 150 kpc for LMC-mass galaxies, and 300 kpc for more massive galaxies. In some cases, host search areas overlap, causing a few candidates to be found in multiple hosts’ search areas. If a candidate is found in more than one searched area, we associate it with the physically closest host, assuming it lies at the same distance. Follow-up observations (see Section VI) will confirm the true hosts of these candidates.
@@
NGC 4244: NGC 4244 (4.20 Mpc) appears to be part of a group of LMC-mass and smaller galaxies, including NGC 4395(4.65 Mpc)). This group of low-mass galaxies has measured TRGB distances between 4 and 5 Mpc (EDD). The galaxies CGCG187-05, UGC 7559, and UGC 7599 lie just beyond NGC 4244’s virial radius, with TRGB distances of 4.85 Mpc, 4.97 Mpc, and 4.72 Mpc, respectively (EDD). These distances place them more than 500 kpc behind NGC 4244, suggesting they belong to the same galaxy group but are unlikely to be its satellites.
NGC 4449: NGC 4449’s (4.16 Mpc) virial radius overlaps with that of M94, which is at a similar distance (4.3 Mpc; EDD), indicating that NGC 4449 and its satellites are likely members of the larger M94 galaxy group.
IC 4182: IC 4182’s (4.24 Mpc) virial radius overlaps with that of M94, which is at a similar distance (4.3 Mpc EDD), indicating that IC 4182 and its satellites are likely members of the larger M94 galaxy group.
NGC 4236: NGC 4236’s (4.31 Mpc) virial radius overlaps with that of UGC 07490; however, the two are unlikely to be associated due to their large difference in velocities. In Cosmicflows-4, NGC 4236 has a measured velocity of 0 km s-1 (EDD), while UGC 07490 has a velocity of 4685 km s-1.
@@
NGC 4395: NGC 4395 (4.65 Mpc) appears to be part of a group of LMC-mass and smaller galaxies, including NGC 4244 (4.20 Mpc). This group has measured TRGB distances between 4 and 5 Mpc (EDD). UGC 07698 and UGC 07605 lie just beyond NGC 4395’s estimated virial radius and appear to belong to the same galaxy group based on their Cosmicflows-4 distances.
NGC 5585: NGC 5585 (6.84 Mpc)) is located on the outskirts of a galaxy group. Its virial radius overlaps that of M101, which is at a similar distance (6.7 Mpc; EDD).
@@
ESO486-G21: For ESO486-G21 (9.7 Mpc), the nearby galaxy NGC 1744 does not have a TRGB distance measurement but has a measured velocity of 741 km s-1 (springbob05). This is similar to ESO486-G21’s velocity of 840 km s-1 (HIPASS). Additionally, multiple Tully-Fischer distance measurements place NGC 1744 within 1 Mpc of ESO486-G21 (e.g. EDD), suggesting that these two galaxies may be associated and part of the same group.